acific Intermountain Utilities Company has a present capital structure (which the company feels is optimal) of 65 percent long-term debt, 5 percent preferred stock, and 30 percent common equity. For the coming year, the company has determined that its optimal capital budget can be externally financed with $55 million of 16 percent first-mortgage bonds sold at par and $20 million of preferred stock costing the company 18 percent. The remainder of the capital budget will be financed with retained earnings. The company’s common stock is presently selling at $15 a share, and next year’s common dividend, D1, is expected to be $2 a share. The company has 22 million common shares outstanding. Next year’s net income available to common stock (including net income from next year’s capital budget) is expected to be $90 million. The company’s past annual growth rate in dividends and earnings has been 7 percent. However, a 5 percent annual growth in earnings and dividends is expected for the foreseeable future. The company’s marginal tax rate is 40 percent. Round your answer to two decimal places. 10.12% would be shown as 10.12. What is the cost of equity? % What is the WACC? %
Answer #1

Cost of equity, rs = D1 / P + g = 2 / 15 + 5% = 18.33%

WACC = wd x rd x (1 - tax) + wps x rps + ws x rs

= 65% x 16% x (1 - 40%) + 5% x 18% + 30% x 18.33%

= 12.64%